Tyrion:

Show all work to multiply quantity 2 plus the square root of negative 25 end quantity times quantity 4 minus the square root of negative 100 end quantity

darkknight:

can you use the eqn tool to write out all that above first?

Tyrion:

1 attachment
darkknight:

do yk foil? first, outer, inner, last

Tyrion:

no

darkknight:

the way it works is like this You have other terms in your problem but ill use a, b, c and d \[(a+b)(c+d)\] outer, multiply a and c, first then a and d then b and c then b and d \[ac+ad+bc+bd\] ill draw this as well

Tyrion:

i thought u find the square root of the numbers first

darkknight:

no need to do that, first foil then simplify, so yk how to foil now?

Tyrion:

ima try it

Tyrion:

aye @darkknight how u do 2 times square root of -25

Tyrion:

would u simplify first

darkknight:

why r u doing 2 time square root of -25? foil you multiply the first terms then the outer, then inner then last

Tyrion:

yes for a times b its 2 and sqaure root -25 right

darkknight:

darkknight wrote:
the way it works is like this You have other terms in your problem but ill use a, b, c and d \[(a+b)(c+d)\] outer, multiply a and c, first then a and d then b and c then b and d \[ac+ad+bc+bd\] ill draw this as well
there is no a times b, you never multiply 2 numbers that are in the same set of parenthesis when expanding

Tyrion:

ohh i read it wrong my bad

darkknight:

ur good

Tyrion:

so it would be 2 times square root -100??

Tyrion:

it says dats 20 i

darkknight:

first 2 times 4, remember foil, the first terms first then yes 2 times sqrt -100, which is 20i, keep g oing

Tyrion:

yuh i did the first one already im juss kurious on wat the i is

Tyrion:

lool

darkknight:

yeah sure, \(i=\sqrt{-1}\) you cant take the square root of a negative number (as a real number) so its imaginary. the sqrt of a negative number isn't a real number, leave it how it is and keep going with the problem, it might cancel later on

Tyrion:

i understand wat u sayin

darkknight:

alr sweet, now move to the inner terms and last

Tyrion:

@snowflake0531 do uk wats next after i did the multiply stuff

snowflake0531:

So do you have (2+ isqrt25)(4- isqrt100)?

Tyrion:

yuh and i did the foil method so wats next

snowflake0531:

Well, write out what you have

snowflake0531:

Also btw you can simplify it to (2+5i)(4-10i)

Tyrion:

AxC=8 AxD=20i BxC=20i BxD=-50

Tyrion:

i thought i was suppose to simplify yesterday dark showed me dis way

snowflake0531:

You're doing it wrong, 2 values are incorrect \[(2+5i)(4-10i) = 8 -20i + 20i + 50\]

Tyrion:

whikh ones

Tyrion:

and dats not a positive 50 its negative

Tyrion:

nor is the first 20i dat was a positive too

snowflake0531:

No, one is positive one is negative

snowflake0531:

-10i(2) and 4(5i)

snowflake0531:

And btw sorry I didn't get the notifs

Tyrion:

im gettin too konfused gettin 2 ways told to me

snowflake0531:

No, we're using the same method, FOIL

Tyrion:

ima juss figure it out on my own

Tyrion:

@darkknight

darkknight:

snowflake0531 wrote:
You're doing it wrong, 2 values are incorrect \[(2+5i)(4-10i) = 8 -20i + 20i + 50\]
read this, I think you may have forgotten the negative sign before the sqrt(-100) in your problem it says 4 minus (sqrt(-100))

Tyrion:

OHHHH i forgot to keep the minus,,so rlly its 8 plus 50??

Tyrion:

kause the 2 insides kankel out

darkknight:

tyrion wrote:
OHHHH i forgot to keep the minus,,so rlly its 8 plus 50??
yea, and you can simplify 8 and 50 also

darkknight:

tyrion wrote:
kause the 2 insides kankel out
correct

Tyrion:

so the answer is 58 or no